# An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors

### v3

@@
Abstract.
Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools. However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity. Large Language Models (LLMs) have been considered promising candidates for addressing those challenges. In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review. Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools. We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance. Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review butå significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 frequently generates responses that are verbose or not compliant with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software post-development issues
1. Introduction
@@
Security defects in software programs can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002). Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022). Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality. During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations. Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository. In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
@@
Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021). Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021). Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013) and weak scalability (Nong et al., 2021).
@@
Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021). However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a). Nong et al. found that even using the training dataset augmented by the latest vulnerability injection technique (Nong et al., 2024b), the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively. Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
@@
Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023). Such models show remarkable capability to understand and generate explanatory natural language responses, thus considered a potential candidate for security code review. We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way. Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently. However, the application of LLMs in the specific field of security code review remains largely unexplored.
@@
Many studies have focused on leveraging LLMs for vulnerability detection. Zhou et al. evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts (Zhou et al., 2024b), while Purba et al. compared LLMs with static analysis tools in detecting software vulnerabilities (Purba et al., 2023). These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks, which was to judge whether there were security defects in the code. They did not require LLMs to provide more detailed information such as the location and type of vulnerability. Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al. (Yang et al., 2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al. (Nong et al., 2024a). However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
@@
Motivated by the above-mentioned limitations, our aim is to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories. Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description and suggested fix. The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review. First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al. (Zhang et al., 2023)). Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. In summary, this study makes the following contributions:
  • We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review.
  • We measured the impact of the randomness of LLMs on the consistency of security code review.
  • We identified quality problems in the responses generated by the top-performing LLM to present its challenges in security code review.
  • We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
@@
Paper Organization. Section 2 surveys the related work. Section 3 describes the methodology employed in this study. Section 4 presents the results, followed by a discussion of the implications of our study results in Section 5. Section 6 clarifies the threats to validity. Finally, Section 7 concludes this work and outlines future directions. The replication package of this study has been provided online (Yu et al., 2024).
2. Related Work
Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, in this section, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or repair.
Charoenwet et al. conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of vulnerability-contributing commits (VCCs) (Charoenwet et al., [n. d.]). By combining several tools, warnings could be produced for 78% of VCCs. However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review. Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgement under various prompts (Zhou et al., 2024b). They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art security defect detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools (Purba et al., 2023). Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description and suggestions for defect remediation.
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers(Bakhshandeh et al., 2023). Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin et al. evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description (Yin and Ni, 2024), found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs. Specifically, in defect localization and description tasks, LLMs exhibited limited overall accuracy and varied performance across different CWE types. These studies have all been limited to quantitatively evaluating LLM performance. In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to gain a comprehensive understanding of the potential and challenges LLMs face in security code review.
Employing specific prompting frameworks to enhance the performances of LLMs in identifying security defects has also been explored. Yang et al. introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection (Yang et al., 2024). Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024a). The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours. These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements. Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
It is worth noting that the aforementioned studies focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs. Inspired by these related works, we aim to leverage the potential of LLMs in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
3. Methodology
3.1. Research Questions
@@
The goal of this work is to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code. To achieve this goal, we divided our study into three Research Questions (RQs), as shown below. The research procedure for each RQ is illustrated in Fig. 1.
@@
RQ1.4: What is the impact of the randomness of LLMs on the consistency of security code review?
Motivation. Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality. It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects. Additionally, due to the inherent randomness of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt. However, this aspect of LLMs has not yet been studied when applied to security tasks. Hence, we conducted a quantitative analysis of the consistency of LLM responses to explore the impact of LLM randomness on the task of security code review.
RQ2: What are the quality problems in LLM-generated responses during security code review?
Motivation. Kabir et al. noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), negatively impacting the quality of the answers. However, for security code review, the quality problems present in responses generated by LLMs have not been thoroughly analyzed in previous studies. Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
@@
Figure 1. An overview of the research procedure for investigating the three RQs
3.2. Data Collection
We leveraged the code review dataset constructed by Yu et al. (Yu et al., 2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (primarily written in Python), and Qt Base and Creator (primarily written in C++). These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where the reviewer commented to identify a security defect. The 614 security-related review comments were reorganized as follows:
First, we filtered out comments published at the patchset level, as such data cannot be located in a specific file to construct a prompt. Then, for each code file to which these comments correspond, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item. Next, we utilized the Gerrit API to retrieve the content of each code file and the commit messages of the patchsets to which they belong. During this procedure, if we cannot obtain the content of the code files through the API, we ignore those files. Lastly, since 97% files in this dataset are Python, C and C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.

### v4

# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

@@
Abstract
Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools. However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity. Large Language Models (LLMs) have been considered promising candidates for addressing those challenges. In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review. Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools. We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance. Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review but significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 frequently generates verbose or non-compliant responses with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
keywords:
@@ Abstract
1 Introduction
@@
Security defects are potential security risks or weaknesses in the system introduced during implementation (Paul et al., 2021). Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002). Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022). Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality. During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations. Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository. In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
@@
Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021). Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021). Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013), and weak scalability (Nong et al., 2021).
@@
Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021). However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a). Nong et al. (2024b) found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively. Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
@@
Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023). Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review. We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way. Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently. However, applying LLMs in the specific field of security code review remains largely unexplored.
@@
Many studies have focused on leveraging LLMs for vulnerability detection. Zhou et al. (2024b) evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while Purba et al. (2023) compared LLMs with static analysis tools in detecting software vulnerabilities. These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code. They did not require LLMs to provide more detailed information, such as the location and type of vulnerability. Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al. (2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al. (2024a). However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
@@
Motivated by the limitations mentioned above, we aim to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories. Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description, and suggested fix. The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review. First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al. (2023)). Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. In summary, this study makes the following contributions:
(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review. (2) We measured the impact of the non-determinism of LLMs on the consistency of security code review. (3) We identified quality problems and quantified hallucinations in the responses generated by the top-performing LLM, aiming to highlight the challenges faced by LLMs in security code review. (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
@@
Paper Organization. Section 6 surveys the related work. Section 2 describes the methodology employed in this study. Section 3 presents the results, followed by a discussion of the implications of our study results in Section 4. Section 5 clarifies the threats to validity. Finally, Section 7 concludes this work and outlines future directions. The replication package of this study has been provided online (Yu et al., 2025).
2 Methodology
2.1 Research Questions
@@
This work aims to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code. We divided our study into three Research Questions (RQs), as shown below. The research procedure for each RQ is illustrated in Fig. 1.
@@
RQ1.4: What is the impact of the non-determinism of LLMs on the consistency of security code review?
Motivation. Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality. It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects. Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs. Non-determinism is mainly caused by decoding configurations (Song et al., 2024), with other hypotheses such as floating point, unreliable GPU calculations (Ouyang et al., 2025), and its sparse MoE architecture failing to enforce per-sequence determinism (La Malfa et al., 2024; Puigcerver et al., 2023). Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
RQ2: What quality problems exist in LLM-generated responses during security code review?
RQ2.1: What is the distribution of quality problems in these responses?
RQ2.2: Which of these quality problems are hallucination-related, and what are their proportions?
Motivation. Kabir et al. (2024) noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers. Among these problems, hallucination in particular has been extensively studied  (Liu et al., 2024; Tonmoy et al., 2024). Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (Huang et al., 2025). If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security. Previous studies have not yet thoroughly examined quality problems present in LLM-generated responses during security code review. RQ2 intends to conduct an in-depth linguistic analysis to investigate possible hallucination issues. First, we investigated the distribution of quality problems in LLMs’ responses, such as Vague statement, Incorrect concept, and Incomplete answer. Then we identified hallucination-related issues among these problems and quantified their proportion to bridge the knowledge gap and highlight potential areas for improvement.
@@
Figure 1: An overview of the research procedure for investigating the three RQs
2.2 Data Collection
We leveraged the code review dataset constructed by Yu et al. (2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python). These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where the reviewer commented to identify a security defect. The 614 security-related review comments were reorganized as follows:
First, we filtered out comments published at the patchset level, as such data cannot be located in a specific file to construct a prompt. Then, for each code file to which these comments correspond, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item. Next, we utilized the Gerrit API to retrieve the content of each code file and the commit messages of the patchsets to which they belong. During this procedure, if we cannot obtain the content of the code files through the API, we ignore those files. Lastly, since 97% of the files in this dataset are Python and C/C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.